OpenStudy (juliebeans):

Can You help me please? My teacher gave me a werid way to do these and i dont understand it I: What is the GCF of 16x^5y^2 + 40x^4y^3? (Points : 4) 8x^4y^2 8x^5y^3 4x^4y^2 4x^5y^3

OpenStudy (juliebeans):

I will metal and fan :)

OpenStudy (anonymous):

Wow...idek

OpenStudy (campbell_st):

ok... so look at the numbers 1st.... what is the GCF of 16 and 40 then look at GCF of\[x^5~~and~~~x^4\] then look at GCF of \[y^2~~~and~~~y^3\]

OpenStudy (juliebeans):

So The Gfc Is 6? Because 3 *2? is 6

OpenStudy (juliebeans):

Hello? @campbell_st

OpenStudy (campbell_st):

6 isn't the GCF of 16 and 40... the factors of 16 are 1, 2, 4, 8, 16 40 1, 2, 4, 5, 8, 10, 20, 40 what is the largest common factor.... of 16 and 40...?

OpenStudy (juliebeans):

8 :P

OpenStudy (juliebeans):

Then We break down everything else next?

OpenStudy (juliebeans):

Hello? ._.

OpenStudy (campbell_st):

ok... now look at \[x^5 = x \times x \times x \times x \times x\] and \[x^4 = x \times x \times x \times x \] what is the largest number of x's that are common to both x^5 and x^4